# ARISE – Adaptive Refinement and Iterative Scenario Engineering

### v1

@@ IV-D Criticality Evaluations
We used SafeBench [37] to confirm compatibility of ARISE with downstream application of the ChatScene. Here, we used an ego vehicle trained with the Soft Actor-Critic (SAC) [18] deep RL methodology provided by SafeBench. The other two trained ego agents, reported in ChatScene, were not provided / published by authors, and therefore were excluded from this evaluation. After an increase of criticality of our scenarios (5 per each of the 8 base scenario categories) using SafeBench, we compared the performance of the ego vehicle on our scenarios across three levels (cf. Table IV) to the ChatScene and its baseline adversary- and knowledge-based scenario generation techniques: Learning-to-collide (LC) [8], AdvSim (AS) [32], Carla Scenario Generator (CS) [6] and Adversarial Trajectory Optimization (AT) [39]. The results show general performance alignment of the ARISE with the baselines, even outperforming them in several metrics (frequency of running red lights (RR), route following stability (RF) and average percentage of route completion (Comp)), while achieving comparable scores on others, reflecting trade-offs in the behavior of the generated scenarios. These metrics are computed within SafeBench, which is configured to deliberately increase scenario criticality for stress-testing; ARISE is focused on automated scenario generation rather than on optimizing this particular downstream task, which could be a focus of a follow-up work. It is important to note that in ChatScene, authors manually polished generated scripts, introducing bias into the generation process, while also increasing criticality of events. Our evaluation showed ChatScene achieved ESR of only 1.4% for automatically generated (no manual involvement) scenarios. In contrast, ARISE operates with minimal human intervention, achieving ESR of 78%. Importantly, in our experiments we found that some of the criticality metrics reported in ChatScene appear unsound and physically unrealistic. For instance, the reported average acceleration (ACC) of 0.705 (cf. Table IV), in combination with virtually guaranteed collision chance, would correspond to implausibly aggressive vehicle dynamics.
Note: The sources will be released upon paper acceptance.
@@ V Future Work and Discussion
As a future enhancement, we plan to integrate the LLM-based evaluator into the TRL as a stricter filter for output scripts. Additionally, allowing users to adjust the weights of scoring criteria would enable finer control, emphasizing specific aspects of script quality. Further improvements could come from experimenting with different LLMs and parameters (e.g., temperature). Since ARISE generates Scenic scripts, future work explores Scenic to OpenSCENARIO translation (to align with the open industry-required formats), expanding input types (e.g., accident reports, non-verbal data), and benchmarking against a wider range of tools. The main objective is to maximize the automation of traffic scenario generation while minimizing human intervention. Accordingly, we focus our downstream evaluation on criticality, primarily to demonstrate that ARISE can also support this stage of the pipeline; however, the impact of generated scenarios on downstream criticality (e.g., SafeBench) is left for future work.
@@ VI Conclusion
We introduce ARISE, an automated tool that converts natural language traffic descriptions into Scenic simulation scripts via an adaptive, feedback-driven process. By incorporating a refinement loop, ARISE effectively translates free-form scenario descriptions into executable code while preserving their semantics. Our experiments show that ARISE achieves significantly higher execution success rates than, while sets a new benchmark, producing both functionally and semantically valid scripts for most inputs. Leveraging a knowledge-augmented LLM, a test-and-repair loop, and improved prompt and knowledge base design, ARISE ensures scripts are syntactically correct and require little to no user intervention. This demonstrates the feasibility of fully automated traffic scenario generation for autonomous driving.

### v2

@@ IV-D Criticality Evaluations
We used SafeBench [37] to confirm compatibility of ARISE with downstream application of the ChatScene. Here, we used an ego vehicle trained with the Soft Actor-Critic (SAC) [18] deep RL methodology provided by SafeBench. The other two trained ego agents, reported in ChatScene, were not provided / published by authors, and therefore were excluded from this evaluation. After an increase of criticality of our scenarios (5 per each of the 8 base scenario categories) using SafeBench, we compared the performance of the ego vehicle on our scenarios across three levels (cf. Table IV) to the ChatScene and its baseline adversary- and knowledge-based scenario generation techniques: Learning-to-collide (LC) [8], AdvSim (AS) [32], Carla Scenario Generator (CS) [6] and Adversarial Trajectory Optimization (AT) [39]. The results show general performance alignment of the ARISE with the baselines, even outperforming them in several metrics (frequency of running red lights (RR), route following stability (RF) and average percentage of route completion (Comp), while achieving comparable scores on others, reflecting trade-offs in the behavior of the generated scenarios. These metrics are computed within SafeBench, which is configured to deliberately increase scenario criticality for stress-testing; ARISE is focused on automated scenario generation rather than on optimizing this particular downstream task, which could be a focus of a follow-up work. It is important to note that in ChatScene, authors manually polished generated scripts, introducing bias into the generation process, while also increasing criticality of events. Our evaluation showed ChatScene achieved ESR of only 1.4% for automatically generated (no manual involvement) scenarios. In contrast, ARISE operates with minimal human intervention, achieving ESR of 78%. Importantly, in our experiments we found that some of the criticality metrics reported in ChatScene appear unrealistic. For instance, the reported average acceleration (ACC) of 0.705 (cf. Table IV), in combination with virtually guaranteed collision chance, would correspond to implausibly aggressive vehicle dynamics.
@@ V Future Work and Discussion
As a future enhancement, we plan to integrate the LLM-based evaluator into the TRL as a stricter filter for output scripts. Additionally, allowing users to adjust the weights of scoring criteria would enable finer control, emphasizing specific aspects of script quality. Further improvements could come from experimenting with different LLMs and parameters (e.g., temperature). Since ARISE generates Scenic scripts, future work will explore Scenic-to-OpenSCENARIO translation (to align with the open industry-required formats), expanding input types (e.g., accident reports, non-verbal data), and benchmarking against a wider range of tools. The main objective is to maximize the automation of traffic scenario generation while minimizing human intervention. Accordingly, we focus our downstream evaluation on criticality, primarily to demonstrate that ARISE can also support this stage of the pipeline; however, the impact of generated scenarios on downstream criticality (e.g., SafeBench) is left for future work.
@@ VI Conclusion
We introduce ARISE, an automated tool that converts natural language traffic descriptions into Scenic simulation scripts via an adaptive, feedback-driven process. By incorporating a refinement loop, ARISE effectively translates free-form scenario descriptions into executable code while preserving their semantics. Our experiments show that ARISE achieves significantly higher execution success rates than the baseline, setting a new benchmark by producing both functionally and semantically valid scripts for most inputs. Leveraging a knowledge-augmented LLM, a test-and-repair loop, and improved prompt and knowledge base design, ARISE ensures scripts are syntactically correct and require little to no user intervention. This demonstrates the feasibility of fully automated traffic scenario generation for autonomous driving.
VII Acknowledgment
This work was supported by the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101076360 and in part by the European Regional Development Fund (ERDF) under grant agreement No. EFRE-AuF-0000866.